OpenStudy (anonymous):

A tennis court has an area of 2808sqft. If it is 6ft longer than twice the width, determine the dimensions of the court?

OpenStudy (saifoo.khan):

Make a simple equation and use the formula of area of rectangle to solve.

OpenStudy (saifoo.khan):

Area = b * h.

OpenStudy (anonymous):

l = 2w + 6 so w(2w+6) = 2808 2w^2 + 6w -2808 = 0 w^2 + 3w - 1404 = 0 (w - 36 )(w + 39 ) = 0 w = 36 ft length = 72 + 6 = 78 ft